OpenStudy (anonymous):

Which is the simplified form of 16 times a to the eighth power all over 12 times a to the fourth power.?

OpenStudy (anonymous):

please help :(

OpenStudy (anonymous):

4a^4/3

OpenStudy (anonymous):

4 over 3. a4 4a2 4a4 4 over 3. a2

OpenStudy (anonymous):

would it be the first one then?

OpenStudy (anonymous):

@jim_thompson5910 hmm hold on.

OpenStudy (anonymous):

i believe so but hold on.

OpenStudy (anonymous):

okayy

jimthompson5910 (jim_thompson5910):

\[\large \frac{16a^8}{12a^4}\] \[\large \frac{16}{12}*\frac{a^8}{a^4}\] \[\large \frac{4}{3}*\frac{a^8}{a^4}\] \[\large \frac{4}{3}*\frac{a^{8-4}}{1}\] \[\large \frac{4}{3}*\frac{a^4}{1}\] \[\large \frac{4}{3}a^4\] so it looks like you're correct

OpenStudy (anonymous):

okay sweet! thanks guys

OpenStudy (anonymous):

Jim is a big help.

jimthompson5910 (jim_thompson5910):

thx str99 and you're welcome